### v1

# ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction

Abstract
The rapid spread of fake news threatens social stability and public trust, rendering its detection an imperative research priority. Although large language models (LLMs) excel at numerous natural language processing tasks with their remarkable contextual understanding and extensive prior knowledge, the time-bounded knowledge coverage and tendency for generating hallucination content reduce their reliability when handling fast-evolving news streams. Furthermore, models trained on existing static datasets also often lack the generalization needed for emerging news topics. To address these challenges, we propose ZoFia, a novel two-stage zero-shot fake news detection framework. First, we introduce Hierarchical Salience to quantify the importance of entities in the news content, and propose the SC-MMR algorithm to effectively select an informative and diverse set of keywords that serve as queries for retrieving up-to-date external evidence. Subsequently, a multi LLM interactive system, in which each agent assumes a distinct role, performs multi-view collaborative analysis and adversarial debate over the news text and its related information, and finally produces an interpretable and robust judgment. Comprehensive experiments on two public datasets demonstrate that ZoFia obviously outperforms existing zero-shot baselines and most of few-shot methods. Our codes will be open-sourced to facilitate related communities.
ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction
Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Tian Wen1, Yuwei Wang1, Min Liu1,2thanks: Corresponding author. 1Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wulvhua24s,jiangxuefeng21b,liumin}@ict.ac.cn
1 Introduction
The rapid spread of fake news through social networks has become prevalent, posing severe threats to key domains such as politics Fisher et al. (2016), economy Bakir and McStay (2018), and livelihood Zhou and Zafarani (2020). The swift advancement of generative models further exacerbates this concern Chen and Shu (2023); Nan et al. (2024); Li et al. (2025, 2024). In this context, developing effective, efficient and interpretable fake news detection methods is imperative.
Fake news detection is formally defined as a binary classification problem. Prior to the rise of Large Language Models (LLMs), related research is predominantly led by training deep models via supervised learning Monti et al. (2019); Kaliyar et al. (2021). LLMs exhibit significant potential in information forensics due to their broad knowledge and strong contextual understanding acquired through pretraining on massive corpora Su et al. (2023). Current research approaches utilizing LLMs for detection generally fall into two categories. The first category employs LLMs as zero-shot reasoners, directly guiding models to judge and explain news authenticity through meticulously designed prompt engineering Jiang et al. (2022) or instruction tuning Pavlyshenko (2023). The second category treats LLMs as auxiliary tools, leveraging their powerful generative capabilities for information augmentation to enhance the performance and robustness of downstream compact classifiers Hu et al. (2024a).
However, directly applying LLMs to fake news detection poses significant challenges. The main problem is that LLMs have a knowledge cutoff Cheng et al. (2024), which means that an LLM’s internal knowledge becomes static after training, limiting its capacity to adapt to dynamic news streams. Furthermore, LLMs are prone to hallucinations Ji et al. (2023), frequently producing plausible but factually incorrect content, which directly undermines its reliability as a misinformation detector. More seriously, existing supervised methods generalize poorly to emerging topics Hoy and Koulouri (2022), and a single LLM used for zero-shot judgment often produces shallow reasoning, leading to superficial analysis.
To address these challenges systematically, we propose ZoFia, a training-free zero-shot framework for fake news detection. ZoFia decouples the complex detection task into two stages. First, it performs entity guided retrieval to integrate the LLM’s internal knowledge with real time external information and builds a comprehensive Multi-Source Information Matrix. Next, an multi-LLM interaction system conducts multidimensional analysis and adversarial debate over this matrix to deliver an explainable and robust judgment. The framework aims to compensate for the knowledge limits of LLMs while fully leveraging their inherent abilities of reasoning and interaction.
Our main contributions can be summarized as follows:
  • We propose ZoFia, a zero-shot fake news detection framework that combines entity-guided retrieval with multi-agent collaboration, effectively addressing LLMs’ knowledge cutoff and hallucinations.
  • We propose two novel algorithms, Hierarchical Salience and SC-MMR, to achieve accurate keyword extraction for news.
  • Experiments conducted on two public real-world datasets demonstrate that our proposed framework outperforms existing zero-shot models and most of few-shot models.
  • Code will be open-sourced to ensure the reproduction and facilitate related communities.
Figure 1: Overall architecture of our proposed ZoFia framework.
2 Related Work
2.1 LLM-based Fake News Detection
Using LLMs for fake news detection is a recent research trend, and current studies explores various capabilities of LLM. Hu et al. Hu et al. (2024a) position LLMs as auxiliary tools and propose that while LLMs cannot directly serve as detectors, they can provide multi-perspective guidance signals to smaller models. GenFEND Nan et al. (2024) focuses on information augmentation, which utilizes LLMs to generate diverse comments simulating different user groups to enrich related information. To address LLM knowledge limitations, Liu et al. Liu et al. (2024) combine internal sample retrieval with external real-time information retrieval, alleviating issues of vague and scarce information.
However, these methods seldom deeply explore other capabilities of LLMs, including inherent reasoning and role-playing. Distinct from prior studies, our framework guides LLMs to play different roles, collaboratively analyzing news from multiple expert perspectives and achieving better and more robust zero-shot detection.
2.2 Multi-agent Collaboration
Multi-agent collaboration as a new paradigm for enhancing LLM capabilities remains in its early stages. Its core concept involves integrating multiple agents to strengthen processing capacity for complex tasks. As pioneering work, Chateval Chan et al. (2023) demonstrated through agent team-based evaluation that this approach simultaneously improves robustness and accuracy in natural language generation. COLA Lan et al. (2024) applied this concept to zero-shot stance detection, assigning different roles to LLMs and combining a debate framework to enhance reasoning, yet failed to effectively mitigate hallucination problems. TruEDebate Liu et al. (2025) introduced multi-stage structured debate into fake news detection, but its implementation revealed a tendency toward premature convergence of viewpoints.
Our work uses a role-playing multi-agent paradigm, but also incorporates external knowledge retrieval and independent modular analysis to reduce hallucination and ensure diversity and independence of debate arguments.
3 Entity-Guided Retrieval
The efficiency of LLMs is constrained by the knowledge cutoff from training data Cheng et al. (2024). This limitation becomes especially apparent when processing rapidly changing real-time news streams, as misinformation usually propagates at rates and patterns that outpace the models’ training cycles Yi et al. (2025). Therefore, enhancing LLMs’ knowledge through external information retrieval is necessary.
We design the entity-guided retrieval stage to acquire real-time and reliable external knowledge for subsequent use by LLM experts. The framework consists of 4 sequential modules. The first 3 modules precisely extract and filter out a set of critical entities from the original news text as the keywords. These keywords are then used as queries for the final retrieval module, which retrieves from both Open Web and Wikipedia. This stage provides LLMs with the essential and up-to-date factual grounding and contextual information.
3.1 Entity Extractor
This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text. This process can be expressed as:
$\{(t_{i},e_{i},c_{i})\}_{i=1}^{N}=\mathcal{M}_{\text{BERT-NER}}(T),$ (1)
where $\mathcal{M}$ is the pre-trained model, $T$ is the input news text. $(t_{i},e_{i},c_{i})$ denotes the recognized entity triplet, $t_{i}$ is the entity token, $e_{i}$ is the entity label, and $c_{i}$ is the confidence score for the corresponding label, expressed as the conditional probability $c_{i}=P(e_{i}|t_{i},T;\mathcal{M}_{\text{BERT-NER}})$.
Due to the large number of recognized entities, we aggregate consecutive entity tokens to form new entity units $U_{e}$. Its confidence score $c(U_{e})$ is calculated by averaging the confidence scoresof all tokens in $U_{e}$, expressed as:
$c(U_{e})=\frac{1}{|U_{e}|}\sum_{t_{i}\in U_{e}}c_{i}.$ (2)
We establish a dynamic confidence threshold to select entities with high confidence scores. Starting from an initial confidence score $\lambda_{\text{init}}$, if the number of selected entities fails to meet the predetermined minimum $n_{\text{min}}$, the algorithm iteratively lowers the confidence score by $\Delta\lambda$ and repeats the selection, until at least $n_{\text{min}}$ entities are obtained.
3.2 Salience Scorer
This module aims to accurately quantify the importance of key entities in news texts, i.e. Entity Salience (ES) Dunietz and Gillick (2014). Although the prior study Bullough et al. (2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by the semantic dilution Hou et al. (2021) problem. In lengthy or unevenly distributed texts, the core semantics of key entities are diluted by minor content, resulting in their true importance being severely underestimated.
We propose a novel Hierarchical Salience that avoids a single, coarse evaluation between an entity and the whole text. It decomposes an entity’s overall importance into two orthogonal and multiplicative components: Local Salience $\mathcal{S}_{\text{local}}$ and Global Salience $\mathcal{S}_{\text{global}}$. Local Salience measures the semantic alignment between the entity and its immediate context, and Global Salience measures how much the local context contributes to the text’s main content.
Formally, for a news text $T$ that is an ordered sequence of sentences, consider any entity $U_{i}$ that appears in sentence $s_{j}$. We define its local context as $\mathcal{C}(U_{i})=s_{j-1}\oplus s_{j}\oplus s_{j+1}$. Aligned with Bullough et al. (2024), we use a unified Sentence-BERT (SBERT) Reimers and Gurevych (2019) encoder $\mathcal{M}_{\text{SBERT}}(\cdot)$ to embed the entity $U_{i}$, its local context $\mathcal{C}(U_{i})$, and the full text $T$ to vectors $\mathbf{v}_{U_{i}}$, $\mathbf{v}_{\mathcal{C}(U_{i})}$, and $\mathbf{v}_{T}$. The hierarchical salience $\mathcal{S}_{\text{hier}}(U_{i})$ is derived from the product of these two components as
$\mathcal{S}_{\text{hier}}(U_{i})=\underbrace{\frac{\mathbf{v}_{U_{i}}\cdot\mathbf{v}_{\mathcal{C}(U_{i})}}{\|\mathbf{v}_{U_{i}}\|\cdot\|\mathbf{v}_{\mathcal{C}(U_{i})}\|}}_{\mathcal{S}_{\text{local}}(U_{i}\mid\mathcal{C}(U_{i}))}\cdot\underbrace{\frac{\mathbf{v}_{\mathcal{C}(U_{i})}\cdot\mathbf{v}_{T}}{\|\mathbf{v}_{\mathcal{C}(U_{i})}\|\cdot\|\mathbf{v}_{T}\|}}_{\mathcal{S}_{\text{global}}(\mathcal{C}(U_{i})\mid T)}.$ (3)
Hierarchical Salience provides a finer and more robust estimate of each entity’s importance to the news content. It serves as the key criterion for entity filtering in subsequent modules.
3.3 Keyword Selector
Input: $\mathcal{D}:\{U_{i}\to\mathcal{S}_{\text{Hier}}(U_{i})\}$: Mapping from entity to its salience score;
$\mathcal{M}_{\text{SBERT}}$: SBERT encoder for entity embeddings;
$\gamma$: Decay factor for termination criterion
Output: $\mathcal{U}_{\text{selected}}$: Final set of selected entities
$\mathcal{U}\leftarrow\text{keys}(\mathcal{D})$; $\mathbf{V}\leftarrow\{\mathcal{M}_{\text{SBERT}}(U_{i})\mid\forall U_{i}\in\mathcal{U}\}$
Compute pairwise similarity matrix $\mathbf{M}_{\text{sim}}$ from $\mathbf{V}$
if $\mathcal{U}_{\text{candidate}}=\emptyset$ then
 return $\mathcal{U}_{\text{selected}}$
$U^{*}\leftarrow\mathop{\arg\max}_{U_{i}\in\mathcal{U}_{\text{candidate}}}\mathcal{D}[U_{i}]$
$\mathcal{U}_{\text{selected}}\leftarrow\{U^{*}\}$; $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}\setminus\{U^{*}\}$
$\text{MMR}_{\text{prev}}^{*}\leftarrow 1.0$
while $\mathcal{U}_{\text{candidate}}\neq\emptyset$ do
 $k\leftarrow|\mathcal{U}_{\text{selected}}|$
 $\lambda_{k}\leftarrow\max(1.0\,,\,1.2-e^{0.2k-1.5}/e)$
 $\text{MMR}_{\text{best}}\leftarrow$-$\infty$; $U_{\text{best}}\leftarrow\text{null}$
 foreach $U_{i}\in\mathcal{U}_{\text{candidate}}$ do
    $\text{sim}_{\max}\leftarrow\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathbf{M}_{\text{sim}}[i,j]$
    $\text{MMR}_{\text{curr}}\leftarrow\lambda_{k}\cdot\mathcal{D}[U_{i}]-(1-\lambda_{k})\cdot\text{sim}_{\max}$
    if $\text{MMR}_{\text{curr}}>\text{MMR}_{\text{best}}$ then
       $\text{MMR}_{\text{best}}\leftarrow\text{MMR}_{\text{curr}}$; $U_{\text{best}}\leftarrow U_{i}$
 if $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}_{\text{prev}}^{*}$ then
    break
 $\text{MMR}_{\text{prev}}^{*}\leftarrow\text{MMR}_{\text{best}}$
 $\mathcal{U}_{\text{selected}}\leftarrow\mathcal{U}_{\text{selected}}\cup\{U_{\text{best}}\}$
 $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}_{\text{candidate}}\setminus\{U_{\text{best}}\}$
return $\mathcal{U}_{\text{selected}}$
Algorithm 1 Salience-Calibrated Maximal Marginal Relevance (SC-MMR)
This module aims to select an informative subset of keywords from the candidate entity set $\mathcal{U}_{\text{selected}}$ to mitigate the query drift Carpineto and Romano (2012) problem. However, this process faces two practical challenges. First, The hierarchical salience score $\mathcal{S}_{\text{hier}}(U_{i})$ is highly sensitive to context granularity, so a fixed screening threshold is not effective. Second, coreference in news introduces high–scoring entities with repeated semantics, which harms the diversity of the keyword set.
To optimize relevance and diversity under these constraints, We propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR) as shown in Algorithm 1. At the $k$-th iteration, SC-MMR evaluates the score of a candidate entity $U_{i}$ by
$\text{MMR}(U_{i})=\lambda_{k}\mathcal{S}_{\text{hier}}(U_{i})-(1-\lambda_{k})\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathcal{S}(U_{i},U_{j}),$ (4a)
$\mathcal{S}(U_{i},U_{j})=\frac{\mathbf{v}_{U_{i}}\cdot\mathbf{v}_{U_{j}}}{\|\mathbf{v}_{U_{i}}\|\,\|\mathbf{v}_{U_{j}}\|}.$ (4b)
The core innovation of SC-MMR is a dynamic decaying weight $\lambda_{k}$. At the early stage, the weight places more emphasis on an entity’s own salience. As the number of selected entities $k$ increases, the weight adaptively decreases for high diversity and low redundancy. The computation of $\lambda_{k}$ is
$\lambda_{k}=\max\left[0.1\,,\,1.2-\frac{\exp(0.2k-0.5)}{e}\right].$ (5)
We further replace termination strategies that rely on a fixed score with a dynamic criterion based on the relative change of the MMR score. The iteration continues only when the next best candidate $\mathcal{U}_{k+1}^{*}$ satisfies $\text{MMR}(\mathcal{U}_{k+1}^{*})>\gamma\cdot\text{MMR}(\mathcal{U}_{k}^{*})$, where $\gamma$ is a decay factor. It lets the algorithm determine the stopping point according to score distribution and prevents a decline in the quality of entities caused by diminishing marginal utility.
3.4 Information Retrieval
This module uses the keyword set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$ distilled in previous modules to build a comprehensive external knowledge base $\mathcal{E}$, serving as supplementary context for the subsequent detection stage. We implement a dual source retrieval that gathers information from the open web and Wikipedia in parallel, integrating the timeliness of the former with the authoritativeness of the latter.
Open Web. We compose all keywords into an aggregated query for the open web $Q_{\text{web}}$ with the following logic:
$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)\land\neg\texttt{news}\land\neg\texttt{wikipedia}.$ (6)
The operator $\land$ requires that results relate to all keywords, which enables strict cross-keyword verification. We explicitly exclude news and wikipedia sources. The former avoids retrieving the duplicate reports to prevent source contamination Deng et al. (2023), and the latter prevents redundancy with the subsequent Wikipedia retrieval. For each returned page, we only extract its summary and search snippet as the raw corpus.
Wikipedia. The summary section of a Wikipedia entry usually provides the most precise definition for an entity. However, a single keyword $k_{i}$ often corresponds to multiple Wikipedia entries, which introduces semantic ambiguity. We build a context-aware disambiguation mechanism that uses the original local context $\mathcal{C}(U_{i})$ of keyword $k_{i}$ in the news text to perform accurate matching.
When Wikipedia returns a list with $M$ candidate senses $\mathcal{O}(k_{i})=\{o_{1},o_{2},\ldots,o_{M}\}$, the mechanism examines each candidate $o_{m}$. It constructs a temporary modified context $\mathcal{C}(U_{i}\leftarrow o_{m})$ by replacing the original entity $U_{i}$ with the description text of $o_{m}$. The optimal sense $o_{i}^{*}$ of $U_{i}$ is defined as the option that maximizes the cosine similarity between the vector of the original context and the vector of the modified context:
$o_{i}^{*}=\mathop{\arg\max}_{o_{m}\in\mathcal{O}(k_{i})}\frac{\mathbf{v}_{\mathcal{C}(U_{i})}\cdot\mathbf{v}_{\mathcal{C}(U_{i}\leftarrow o_{m})}}{\|\mathbf{v}_{\mathcal{C}(U_{i})}\|\|\mathbf{v}_{\mathcal{C}(U_{i}\leftarrow o_{m})}\|}.$ (7)
where $\mathbf{v}$ denotes the context vectors embedded by the pretrained SBERT $\mathcal{M}_{\text{SBERT}}$. After identifying the unique entry, we extract the first 3 sentences of its summary as supplementary material.
4 Multi-LLM Interaction
Entity-guided retrieval supplies the framework with timely external information, which together with the prior knowledge of LLMs builds a Multi-Source Information Matrix. This stage employs a dual-state multi-LLM interaction system that fully exploits this matrix to mitigate hallucinations Feng et al. (2024) and ensure accurate detection.
This stage operates in two orthogonal interaction states. LLM Collaboration performs parallel analyses across multiple agents to reduce inferential variance. LLM Debate and Judgment introduces a adversarial debate to reduce systemic bias. These two states form a complete reasoning chain from mining divergent evidences to making a convergent judgment.
4.1 LLM Collaboration
LLM Collaboration state aims to reduce inferential variance. Through parallel analysis by multiple agents, it transforms Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate. As illustrated in Figure 2, the matrix systematically integrates 4 information sources that are orthogonal and complementary in evidence distribution:
  • In-news Information: It comes from the original text of the target news and provides the basic core content.
  • Out-of-news Information: It comes from Open Web retrieval and offers the broadest instant facts.
  • LLM internal Knowledge: It comes from the model’s prior knowledge and provides generalized common sense.
  • LLM external Knowledge: It comes from Wikipedia retrieval and provides the most precise summary for the core entities.
The subsequent analysis decomposes the tasks and assigns them to agents with distinct roles, so that each agent focuses on a specific quadrant of the matrix for specialized processing.
Figure 2: The diagram of Multi-Source Information Matrix and the quadrants used by LLM agents.
4.1.1 Linguist
Following prior studies Shahid et al. (2022)Zhou and Zafarani (2018), the linguist agent is designed to systematically divide the text into 5 linguistic dimensions that are strongly associated with misinformation:
  • Sentence: Lexical complexity, sentence length, and formality of tone.
  • Word: Frequency of superlatives, affective language, and pronoun distribution.
  • Grammar: Patterns of reported speech, passive voice, and negation.
  • Emotion: Proportion of affective terms in the text and the headline, and the degree of incendiary tone.
  • Information Quality: Presence of clickbait, information overload, or context mismatch.
To maintain objective independence, each dimension is evaluated in an isolated session. For each dimension, LLM explicitly indicates whether it reflects the news is real or fake. This agent uses 2 quadrants of Multi-source Information Matrix: LLM internal knowledge and in-news information from the original news text.
4.1.2 Domain-Specific Expert
It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as “economist” or “journalist”.
The expert with this role then analyses along the following 2 dimensions:
  • Knowledge Concordance: Examine all claims, viewpoints, and details for sound reasoning; identify departures from common sense.
  • Logical Integrity: Examine argument-to-conclusion coherence with domain-specific common sense; identify logical errors or unsupported leaps.
This agent uses 3 quadrants of Multi-source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge from Wikipedia.
4.1.3 Claim Verification
Existing research Niu et al. (2024) demonstrates that claim-based fact-checking effectively serves as a reference for LLM-based detection. In our system, a serial pipeline composed of Claim Extractor and Claim Verifier deconstructs and verifies factual claims in the news text. These 2 agents use 3 quadrants of Multi-source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
Claim Extractor. $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
$\mathcal{M}_{\text{extractor}}(T)\rightarrow\{q_{\text{core}},\{q_{\text{sub}_{1}},\dots,q_{\text{sub}_{m}}\}\}$ (8)
where $q_{\text{core}}$ is the core claim that determines the veracity of the news, and $q_{\text{sub}}$} is a collection of supporting subclaims. All outputs are restated as concise and objective declarative sentences.
Claim Verifier. This agent verifies each claim $q$ independently. To ensure precision and control cost Purwar et al. (2024), a simple retrieval augmented generation (RAG) Lewis et al. (2020) is implemented to build a highly relevant context $\mathcal{C}_{\text{rel}}(q)$ for each $q$ from the Open Web corpus $\mathcal{I}_{\text{web}}$.
The context consists of text chunks $c_{j}$ whose cosine similarity with the claim representation $\mathbf{v}_{q}$ exceeds a threshold $\theta_{\text{sim}}$:
$\mathcal{C}_{\text{rel}}(q)=\{c_{j}\in\text{top-k}(\mathcal{I}_{\text{web}},q)\mid\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c_{j}}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c_{j}}\|}\geq\theta_{\text{sim}}\}.$ (9)
All decisions are strictly based on $\mathcal{C}_{\text{rel}}(q)$. The final output includes a clear label (“Supports”, “Refutes”, or “Not Enough Information”), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ for suppressing hallucinations.
4.2 LLM Debate and Judgment
LLM Debate and Judgment state aims to suppress systemic bias. We introduce a multi-round adversarial debate framework that forces LLMs to explore both supporting and refuting views of the truthfulness of news equally. This design directly mitigates the thought degeneration (DoT) Liang et al. (2023) phenomenon that often appears in a single linear reasoning chain.
Input: $\mathbb{E}$: The complete evidence pool
Output: $D$: The final decision $\{\textit{Real},\textit{Fake}\}$
$\mathbb{H}\leftarrow\emptyset$; $A_{\text{con}}\leftarrow\text{null}$; $D\leftarrow\textit{Insufficient}$
while $D=\text{Insufficient}$ do
 $A_{\text{pro}}\leftarrow\mathcal{M}_{\text{pro}}.\text{GenerateArgument}(\mathbb{E},A_{\text{con}})$
 $A_{\text{con}}\leftarrow\mathcal{M}_{\text{con}}.\text{GenerateArgument}(\mathbb{E},A_{\text{pro}})$
 $\mathbb{H}\leftarrow\mathbb{H}\cup\{(A_{\text{pro}},A_{\text{con}})\}$
 $D\leftarrow\text{Judge.Assess}(\mathbb{H})$
return $D$
Algorithm 2 Dynamic Adversarial Debate
As shown in Algorithm 2, a pair of debate agents $\mathcal{M}_{\text{pro}}$ and $\mathcal{M}_{\text{con}}$ act as opposing reasoners based on the evidence pool $\mathbb{E}$. In each round, the active debater first rebuts the opponent’s last argument and then presents a new argument. After each exchange, a judge agent evaluates the debate history $\mathcal{H}$ and outputs a ternary judgment $D$.
This dynamic termination mechanism ensures that the debate stops once the information is sufficient for decision, effectively balancing the depth and efficiency of reasoning. The debate history $\mathbb{H}$ provides a traceable reasoning chain composed of pro. and con. arguments, making the final judgment highly interpretable.
5 Experiment
5.1 Experimental Setting
Herein we discuss main experimental settings.
Datasets. Following previous state-of-the-art work Liu et al. (2024), our experiments are conducted on two widely recognized fake news datasets: GossipCop and PolitiFact Shu et al. (2020). GossipCop focuses on entertainment, mainly Hollywood celebrity news; PolitiFact focuses on politics, drawing on fact-checks of U.S. political figures. Considering that some links in the original datasets have become invalid, we adopt subsets publicly released in Su et al. (2023).
Metrics and Baselines. We use the accuracy (Acc.) and F1-score (F1) as the main metrics. To reflect performance on detecting fake news, we compute the F1-Score for the fake class. ZoFia compares against two groups of baselines. The first group includes 5 advanced Few-Shot methods: MDFEND Nan et al. (2021), ARG Hu et al. (2024b), PSM Ni et al. (2020), DKFND Liu et al. (2024), and KPL Jiang et al. (2022), with metrics retrieved from the existing literature Jin et al. (2024); Hu et al. (2024c); Liu et al. (2024). The second group consists of 3 representative Zero-Shot methods that we implement for comparison: Standard Prompt, Auto-CoT Zhang et al. (2022), and ReAct Yao et al. (2023) equipped with a search API.
Implementation details. We set the dynamic threshold for entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor for MMR as $\gamma=0.5$. The maximum entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry. The minimum similarity threshold for claim extraction is $\theta_{\text{sim}}=0.1$. The NER model is selected as dslim/bert-base-NER Tjong Kim Sang and De Meulder (2003), and the SBERT model is selected as all-MiniLM-L6-v2 Reimers and Gurevych (2019). All zero-shot methods we implement use DeepSeek-V3 DeepSeek‑AI (2024) with temperature set to 0.1 as the base model, and Brave API [^1] https://brave.com/search/api/ as the search API.
5.2 Main Experiment
The experimental results are presented in Table 1. ZoFia demonstrates exceptional performance on both datasets, consistently outperforming existing zero-shot and few-shot baselines. On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.9%) and F1-score (89.8%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin, proving its powerful efficacy in fact-checking tasks.
Table 1: Accuracy (Acc.) and F1-score (F1) comparison of few-shot / zero-shot methods on PolitiFact and GossipCop. The bold and underlined denote the best and second-best performance.
Method PolitiFact GossipCop Acc. F1 Acc. F1 MDFEND 65.5 62.3 41.2 40.2 ARG 74.0 67.1 61.4 42.3 PSM 70.0 49.1 77.4 41.7 DKFND 87.0 67.1 82.4 42.3 KPL 58.3 60.4 42.7 42.1 Standard Prompt 79.0 40.2 61.0 28.1 Auto-CoT 89.7 73.7 56.7 43.3 ReAct (Search API) 74.7 67.6 74.0 47.3 ZoFia (Ours) 91.9 89.8 78.5 61.2
The detection task on the GossipCop dataset presents greater challenges, as its content often proves difficult to verify due to subjectivity and factual ambiguity. In this scenario, ZoFia achieves the highest F1-score (61.2%), surpassing all baseline models. Although DKFND attains slightly higher accuracy, its F1-score is notably lower. In contrast, ZoFia’s superior F1-score demonstrates more balanced and robust detection performance across both true and fake news categories while maintaining high precision.
The experimental results clearly indicate that as a zero-shot method, ZoFia consistently outperforms the vast majority of few-shot and zero-shot models, showcasing its advancement and practical utility.
5.3 Ablation Study
To systematically assess the effectiveness of each component in ZoFia, we conduct extensive ablation on the GossipCop dataset.
Figure 3: Performance comparison of ZoFia’s keyword extraction with other extraction methods.
Effectiveness of Entity-Guided Retrieval. As shown in Figure 3, replacing our keyword extraction module with the classic TF-IDF Sparck Jones (1972) algorithm lowers the F1-Score by 9.1%. This finding emphasizes the need for our Hierarchical Salience. We also evaluate Sentence Salience Bullough et al. (2024) strategy, which retrieves a full sentence. The result shows an extreme imbalance: high Accuracy with low F1-Score. The reason is that using sentences as queries tends to retrieve pages that are highly similar to the original news. This source contamination Deng et al. (2023) induces confirmation bias in LLM agents.
Table 2: Results of ablation study on GossipCop
Component Acc. F1 ZoFia 78.5 61.2 w/o Open Web retrieval 71.7 58.1 w/o Wikipedia retrieval 77.7 59.9 w/o Open Web and Wikipedia retrieval 71.3 58.2 w/o linguist analysis 77.8 58.9 w/o expert analysis 76.8 57.3 w/o linguist and expert analyses 73.5 58.9 w/o claim verification 79.2 60.5 w/o debate 75.0 59.2
Component Analysis of Multi-LLM Interaction. As shown in Table 2, removing Open Web retrieval causes a marked decline, which demonstrates the importance of timely external knowledge. Performance decreases further when all external knowledge sources are removed. For LLM collaboration, removing the linguist or expert weakens the framework. The expert contributes more, with a 3.9% drop in F1-Score. Moreover, removing both causes cumulative degradation. Removing the debate reduces the F1-Score by 2.0%, demonstrating its role in suppressing single-view bias.
5.4 Sensitivity Analysis of SC-MMR
The dynamic weight $\lambda_{k}$ is the most important parameter of SC-MMR introduced in Section 3. We conduct a sensitivity analysis before designing it. First, $\lambda_{k}$ is fixed to 0.5, indicating equal emphasis on relevance and diversity in keyword selection. As shown in Figure 4, we observe that when the number of selected keywords $k\leq 6$, F1-Score remains stable at a relatively high level, but once $k>6$, F1-Score drops sharply. This pattern indicates that too many keywords introduce noise and hurt detection performance.
Figure 4: The effect of the number of keywords $k$ on performance (F1-Score).
Based on this finding, we designed $\lambda_{k}$ a decreasing function with an increasing decay rate, as shown in Eq. 5. It remains high for small $k$ to prioritize relevance, reaching approximately 0.45 at $k=6$, but rapidly decaying for $k>6$. This design lowers the MMR score and smoothly applies the relative termination rule. It allows SC-MMR to adapt the size of the keyword set while avoiding performance loss due to redundancy.
6 Conclusion
We propose ZoFia, a training-free zero-shot framework for fake news detection, to address the fundamental conflict between the static prior knowledge of large language models (LLMs) and the uncertainty of dynamic news streams. ZoFia uses a two-stage pipeline to tackle this challenge. In the first stage, our novel Hierarchical Salience and SC-MMR algorithms extract a set of informative and low-redundancy keywords from the news text, and then build a multi-source information matrix that integrates LLM internal and external knowledge with in-news and out-of-news information. In the second stage, a dual-state multi-LLM interaction processes this matrix. The agents first perform collaborative analysis to reduce inferential variance, and then conduct adversarial debate to suppress systemic bias, finally delivering an interpretable and robust judgment. Comprehensive experiments on two widely used datasets show that ZoFia outperforms existing zero-shot baselines and the majority of few-shot methods. Detailed ablation study further demonstrates the importance and effectiveness of each core component in the framework.
Limitations
Though ZoFia show strong detection capacities, its application and evaluation face multiple constraints. Due to copyright and privacy concerns, there has been a recent lack of high-quality, and continuously updated public datasets. It prevents us from evaluating ZoFia on the most recent news. Building the next generation of benchmark datasets that meet ethical standards and reflect real-world information dynamics is a critical step in this field.
From the framework’s perspective, ZoFia depends on multiple LLMs, which introduces nontrivial computation cost and inference latency. The efficiency of using external retrieval can also be improved. Since it is not our main focus, we integrate only a lightweight retrieval augmented generation (RAG) module in claim verification. Future work that adopts more advanced RAG architectures, such as re-ranking model and more advanced mechanism, may further strengthen the exploitation of external knowledge. We plan to conduct more comprehensive experimental benchmark evaluation over the detection capabilities of LLMs and Multi-agent system for fake news Guo et al. (2025); Kuntur et al. (2024); Jiang et al. (2024).
ZoFia currently focuses on the text modality. Modern misinformation increasingly appears in multi-modal form that combines images and text. Extending ZoFia to the multimodal domain has strong potential. One direction is to introduce vision language models (VLMs) as a dedicated visual expert. Another is to study how different modalities interact during the debate process to achieve effective fusion.

### v3

@@
 Abstract
-The rapid spread of fake news threatens social stability and public trust, rendering its detection an imperative research priority.
-Although large language models (LLMs) excel at numerous natural language processing tasks with their remarkable contextual understanding and extensive prior knowledge, the time-bounded knowledge coverage and tendency for generating hallucination content reduce their reliability when handling fast-evolving news streams.
-Furthermore, models trained on existing static datasets also often lack the generalization needed for emerging news topics.
-To address these challenges, we propose ZoFia, a novel two-stage zero-shot fake news detection framework.
-First, we introduce Hierarchical Salience to quantify the importance of entities in the news content, and propose the SC-MMR algorithm to effectively select an informative and diverse set of keywords that serve as queries for retrieving up-to-date external evidence.
-Subsequently, a multi LLM interactive system, in which each agent assumes a distinct role, performs multi-view collaborative analysis and adversarial debate over the news text and its related information, and finally produces an interpretable and robust judgment.
-Comprehensive experiments on two public datasets demonstrate that ZoFia obviously outperforms existing zero-shot baselines and most of few-shot methods.
-Our codes will be open-sourced to facilitate related communities.
+The rapid spread of fake news threatens social stability and public trust, highlighting the urgent need for its effective detection.
+Although large language models (LLMs) show potential in fake news detection, they are limited by knowledge cutoff and easily generate factual hallucinations when handling time-sensitive news.
+Furthermore, the thinking of a single LLM easily falls into early stance locking and confirmation bias, making it hard to handle both content reasoning and fact checking simultaneously.
+To address these challenges, we propose ZoFia, a two-stage zero-shot fake news detection framework.
+In the first retrieval stage, we propose novel Hierarchical Salience and Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to extract core entities accurately, which drive dual-source retrieval to overcome knowledge and evidence gaps.
+In the subsequent stage, a multi-agent system conducts multi-perspective reasoning and verification in parallel and achieves an explainable and robust result via adversarial debate.
+Comprehensive experiments on two public datasets show that ZoFia outperforms existing zero-shot baselines and even most few-shot methods.
+Our code has been open-sourced to facilitate the research community at https://github.com/SakiRinn/ZoFia.
 ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction
-Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Tian Wen1, Yuwei Wang1, Min Liu1,2thanks: Corresponding author.
+Lvhua Wu1,2thanks: Equal contribution., Xuefeng Jiang1,211footnotemark: 1, Sheng Sun1, Yan Lei1,2, Tian Wen1,2, Yuwei Wang1, Min Liu1,2thanks: Corresponding author.
 1Institute of Computing Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wulvhua24s,jiangxuefeng21b,liumin}@ict.ac.cn
@@
 (2024); Li et al.
-(2025, 2024).
-In this context, developing effective, efficient and interpretable fake news detection methods is imperative.
-Fake news detection is formally defined as a binary classification problem.
-Prior to the rise of Large Language Models (LLMs), related research is predominantly led by training deep models via supervised learning Monti et al.
+(2025, 2024a).
+In this context, developing effective, efficient and interpretable fake news detection methods is indispensable.
+Early studies mainly rely on supervised learning Monti et al.
 (2019); Kaliyar et al.
+(2021) to train detection models, but their dependence on large-scale labeled data makes it hard to adapt to emerging news topics Hoy and Koulouri (2022).
+Large language models (LLMs), with broad pretraining knowledge and strong contextual understanding Su et al.
+(2023), significantly advance few-shot and zero-shot methods, providing new opportunities to overcome this bottleneck.
+In few-shot methods, LLMs either serve as auxiliary tools to perform data augmentation for downstream classifiers Hu et al.
+(2024a), or act directly as detectors through prompt learning Jiang et al.
+(2022) and instruction tuning Pavlyshenko (2023), but they still do not fully escape reliance on training data.
+By contrast, zero-shot methods that do not require labeled samples are a more promising paradigm, which mainly guide LLM reasoning by context engineering Zhang and Gao (2023) and agentic architectures Li et al.
+(2024b).
+However, zero-shot methods face reliability challenges.
+On the one hand, the internal knowledge of LLMs is limited by knowledge cutoff Cheng et al.
 (2024).
-LLMs exhibit significant potential in information forensics due to their broad knowledge and strong contextual understanding acquired through pretraining on massive corpora Su et al.
+Without the external information, they easily produce factual hallucinations Ji et al.
+(2023) when handling dynamic news events.
+On the other hand, LLMs tend to lock their stance early Echterhoff et al.
+(2024), where confirmation bias Wan et al.
+(2025) drives subsequent reasoning to merely rationalize the initial judgment Wang et al.
 (2025).
-Current research approaches utilizing LLMs for detection generally fall into two categories.
-The first category employs LLMs as zero-shot reasoners, directly guiding models to judge and explain news authenticity through meticulously designed prompt engineering Jiang et al.
-(2022) or instruction tuning Pavlyshenko (2023).
-The second category treats LLMs as auxiliary tools, leveraging their powerful generative capabilities for information augmentation to enhance the performance and robustness of downstream compact classifiers Hu et al.
-(2024a).
-However, directly applying LLMs to fake news detection poses significant challenges.
-The main problem is that LLMs have a knowledge cutoff Cheng et al.
-(2024), which means that an LLM’s internal knowledge becomes static after training, limiting its capacity to adapt to dynamic news streams.
-Furthermore, LLMs are prone to hallucinations Ji et al.
-(2023), frequently producing plausible but factually incorrect content, which directly undermines its reliability as a misinformation detector.
-More seriously, existing supervised methods generalize poorly to emerging topics Hoy and Koulouri (2022), and a single LLM used for zero-shot judgment often produces shallow reasoning, leading to superficial analysis.
-To address these challenges systematically, we propose ZoFia, a training-free zero-shot framework for fake news detection.
-ZoFia decouples the complex detection task into two stages.
-First, it performs entity guided retrieval to integrate the LLM’s internal knowledge with real time external information and builds a comprehensive Multi-Source Information Matrix.
-Next, an multi-LLM interaction system conducts multidimensional analysis and adversarial debate over this matrix to deliver an explainable and robust judgment.
-The framework aims to compensate for the knowledge limits of LLMs while fully leveraging their inherent abilities of reasoning and interaction.
-Our main contributions can be summarized as follows:
-  • We propose ZoFia, a zero-shot fake news detection framework that combines entity-guided retrieval with multi-agent collaboration, effectively addressing LLMs’ knowledge cutoff and hallucinations.
-  • We propose two novel algorithms, Hierarchical Salience and SC-MMR, to achieve accurate keyword extraction for news.
-  • Experiments conducted on two public real-world datasets demonstrate that our proposed framework outperforms existing zero-shot models and most of few-shot models.
-  • Code will be open-sourced to ensure the reproduction and facilitate related communities.
+Consequently, when external retrieval is introduced into a single LLM, it easily takes a cognitive shortcut that substitutes whether information can be retrieved for factual veracity, which seriously weakens content reasoning on the original news.
+This effect becomes much stronger when the retrieved evidence is irrelevant or contradictory Tan et al.
+(2024).
+Therefore, we observe that the inherent reasoning defects of a single LLM prevent it from performing both news content reasoning and external evidence verification well simultaneously.
+We argue that content reasoning and fact checking should be decoupled and reliable judgment can be achieved through multi-agent interaction.
+Based on this motivation, we propose ZoFia, a two-stage zero-shot framework for fake news detection.
+In the first entity-guided retrieval stage, we introduce Hierarchical Salience to address semantic dilution Hou et al.
+(2021) of news entities, which fully leverages global and local semantics to score entity salience, and design a novel Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to accurately extract core entities.
+These entities then drive dual-source retrieval from Wikipedia and Open Web to mitigate knowledge cutoff and effectively suppress hallucinations.
+In the subsequent multi-LLM interaction stage, ZoFia assigns agents to perform content reasoning and fact checking in parallel and breaks stance locking of a single LLM via adversarial debate, ensuring the robustness and interpretability of the final verdict.
+To sum up, our main contributions are outlined as follows:
+  • We propose ZoFia, a retrieval-augmented multi-agent zero-shot framework for fake news detection that effectively overcomes the inherent reasoning flaws of a single LLM.
+  • We introduce a novel granularity-aware metric Hierarchical Salience and design SC-MMR algorithm to extract news entities accurately for efficient retrieval.
+  • Comprehensive experiments on two public datasets demonstrate that ZoFia outperforms existing zero-shot baselines and even most few-shot methods.
+  • We open-source our code to facilitate the research community at https://github.com/SakiRinn/ZoFia.
 Figure 1: Overall architecture of our proposed ZoFia framework.
 2 Related Work
-2.1 LLM-based Fake News Detection
-Using LLMs for fake news detection is a recent research trend, and current studies explores various capabilities of LLM.
-Hu et al.
-Hu et al.
-(2024a) position LLMs as auxiliary tools and propose that while LLMs cannot directly serve as detectors, they can provide multi-perspective guidance signals to smaller models.
-GenFEND Nan et al.
-(2024) focuses on information augmentation, which utilizes LLMs to generate diverse comments simulating different user groups to enrich related information.
-To address LLM knowledge limitations, Liu et al.
-Liu et al.
-(2024) combine internal sample retrieval with external real-time information retrieval, alleviating issues of vague and scarce information.
-However, these methods seldom deeply explore other capabilities of LLMs, including inherent reasoning and role-playing.
-Distinct from prior studies, our framework guides LLMs to play different roles, collaboratively analyzing news from multiple expert perspectives and achieving better and more robust zero-shot detection.
-2.2 Multi-agent Collaboration
-Multi-agent collaboration as a new paradigm for enhancing LLM capabilities remains in its early stages.
-Its core concept involves integrating multiple agents to strengthen processing capacity for complex tasks.
-As pioneering work, Chateval Chan et al.
-(2023) demonstrated through agent team-based evaluation that this approach simultaneously improves robustness and accuracy in natural language generation.
-COLA Lan et al.
-(2024) applied this concept to zero-shot stance detection, assigning different roles to LLMs and combining a debate framework to enhance reasoning, yet failed to effectively mitigate hallucination problems.
+LLM in Fake News Detection.
+The application of Large Language Models (LLMs) has become a research frontier in fake news detection, primarily divided into few-shot and zero-shot paradigms.
+In few-shot methods, researchers often use LLMs as auxiliary tools for data augmentation Hu et al.
+(2024a); Nan et al.
+(2024), or as detectors via post-training Jiang et al.
+(2022); Pavlyshenko (2023).
+However, these methods fail to completely eliminate dependence on labeled data.
+Zero-shot methods directly guide model reasoning through context engineering Zhang and Gao (2023) and agentic architectures Li et al.
+(2024b); Liu et al.
+(2024), making judgments without labeled samples.
+Nevertheless, these methods fail to address the confirmation bias arising from a single reasoning chain.
+We decouple retrieval from reasoning via a multi-role multi-agent system and finally aggregate all evidence for judgment, achieving more comprehensive and robust zero-shot discrimination.
+Multi-Agent System.
+Multi-Agent Systems (MAS) have emerged as an effective paradigm to enhance LLMs for complex tasks.
+The pioneering work Chateval Chan et al.
+(2023) demonstrates that MAS improves both the robustness and accuracy of generation tasks.
+Subsequent studies introduce this paradigm to reasoning tasks.
+For instance, COLA Lan et al.
+(2024) designs a collaboration framework for stance detection, but it remains limited to analyzing the original text.
 TruEDebate Liu et al.
-(2025) introduced multi-stage structured debate into fake news detection, but its implementation revealed a tendency toward premature convergence of viewpoints.
-Our work uses a role-playing multi-agent paradigm, but also incorporates external knowledge retrieval and independent modular analysis to reduce hallucination and ensure diversity and independence of debate arguments.
-3 Entity-Guided Retrieval
-The efficiency of LLMs is constrained by the knowledge cutoff from training data Cheng et al.
-(2024).
-This limitation becomes especially apparent when processing rapidly changing real-time news streams, as misinformation usually propagates at rates and patterns that outpace the models’ training cycles Yi et al.
-(2025).
-Therefore, enhancing LLMs’ knowledge through external information retrieval is necessary.
-We design the entity-guided retrieval stage to acquire real-time and reliable external knowledge for subsequent use by LLM experts.
-The framework consists of 4 sequential modules.
-The first 3 modules precisely extract and filter out a set of critical entities from the original news text as the keywords.
-These keywords are then used as queries for the final retrieval module, which retrieves from both Open Web and Wikipedia.
-This stage provides LLMs with the essential and up-to-date factual grounding and contextual information.
+(2025) applies structured debate to fake news detection, but its implementation tends to cause premature consensus convergence.
+Recent agentic systems Zhang et al.
+(2026) also integrate reasoning more tightly with external knowledge access.
+Our ZoFia introduces external information retrieval and independent modular analysis.
+This design aims to fundamentally mitigate knowledge cutoff and ensure the diversity and independence of arguments and analyses.
+3 Stage 1: Entity-Guided Retrieval
+This stage aims to acquire reliable external knowledge and instant factual evidence for the subsequent multi-agent system.
+It consists of four sequential modules.
+The first three modules precisely extract a set of core entities from the original news.
+These entities serve as keywords and are concatenated into a query for the final retrieval module, which retrieves from Open Web and Wikipedia.
 3.1 Entity Extractor
-This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text. This process can be expressed as:
+This module first uses a pre-trained BERT-NER Tjong Kim Sang and De Meulder (2003) model to perform named entity recognition (NER) on the news text.
+We adopt a lightweight BERT-NER model to obtain stable span-level confidence scores for dynamic filtering.
+This process can be expressed as:
 $\{(t_{i},e_{i},c_{i})\}_{i=1}^{N}=\mathcal{M}_{\text{BERT-NER}}(T),$ (1)
@@
 3.2 Salience Scorer
-This module aims to accurately quantify the importance of key entities in news texts, i.e.
-Entity Salience (ES) Dunietz and Gillick (2014).
+This module aims to accurately quantify the importance of news entities, namely Entity Salience (ES) Dunietz and Gillick (2014).
 Although the prior study Bullough et al.
-(2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by the semantic dilution Hou et al.
-(2021) problem.
-In lengthy or unevenly distributed texts, the core semantics of key entities are diluted by minor content, resulting in their true importance being severely underestimated.
+(2024) has demonstrated that bi-encoder architectures can efficiently estimate entity salience, their performance is often limited by semantic dilution Hou et al.
+(2021) problem, which leads to severe underestimation of the importance of key entities.
 We propose a novel Hierarchical Salience that avoids a single, coarse evaluation between an entity and the whole text.
@@
 3.3 Keyword Selector
-Input: $\mathcal{D}:\{U_{i}\to\mathcal{S}_{\text{Hier}}(U_{i})\}$: Mapping from entity to its salience score;
-$\mathcal{M}_{\text{SBERT}}$: SBERT encoder for entity embeddings;
-$\gamma$: Decay factor for termination criterion
-Output: $\mathcal{U}_{\text{selected}}$: Final set of selected entities
-$\mathcal{U}\leftarrow\text{keys}(\mathcal{D})$; $\mathbf{V}\leftarrow\{\mathcal{M}_{\text{SBERT}}(U_{i})\mid\forall U_{i}\in\mathcal{U}\}$
-Compute pairwise similarity matrix $\mathbf{M}_{\text{sim}}$ from $\mathbf{V}$
-if $\mathcal{U}_{\text{candidate}}=\emptyset$ then
- return $\mathcal{U}_{\text{selected}}$
-$U^{*}\leftarrow\mathop{\arg\max}_{U_{i}\in\mathcal{U}_{\text{candidate}}}\mathcal{D}[U_{i}]$
-$\mathcal{U}_{\text{selected}}\leftarrow\{U^{*}\}$; $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}\setminus\{U^{*}\}$
-$\text{MMR}_{\text{prev}}^{*}\leftarrow 1.0$
-while $\mathcal{U}_{\text{candidate}}\neq\emptyset$ do
- $k\leftarrow|\mathcal{U}_{\text{selected}}|$
- $\lambda_{k}\leftarrow\max(1.0\,,\,1.2-e^{0.2k-1.5}/e)$
- $\text{MMR}_{\text{best}}\leftarrow$-$\infty$; $U_{\text{best}}\leftarrow\text{null}$
- foreach $U_{i}\in\mathcal{U}_{\text{candidate}}$ do
-    $\text{sim}_{\max}\leftarrow\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathbf{M}_{\text{sim}}[i,j]$
-    $\text{MMR}_{\text{curr}}\leftarrow\lambda_{k}\cdot\mathcal{D}[U_{i}]-(1-\lambda_{k})\cdot\text{sim}_{\max}$
-    if $\text{MMR}_{\text{curr}}>\text{MMR}_{\text{best}}$ then
-       $\text{MMR}_{\text{best}}\leftarrow\text{MMR}_{\text{curr}}$; $U_{\text{best}}\leftarrow U_{i}$
- if $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}_{\text{prev}}^{*}$ then
-    break
- $\text{MMR}_{\text{prev}}^{*}\leftarrow\text{MMR}_{\text{best}}$
- $\mathcal{U}_{\text{selected}}\leftarrow\mathcal{U}_{\text{selected}}\cup\{U_{\text{best}}\}$
- $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}_{\text{candidate}}\setminus\{U_{\text{best}}\}$
-return $\mathcal{U}_{\text{selected}}$
-Algorithm 1 Salience-Calibrated Maximal Marginal Relevance (SC-MMR)
 This module aims to select an informative subset of keywords from the candidate entity set $\mathcal{U}_{\text{selected}}$ to mitigate the query drift Carpineto and Romano (2012) problem.
@@
 First, the hierarchical salience score $\mathcal{S}_{\text{hier}}(U_{i})$ is highly sensitive to context granularity, so a fixed screening threshold is not effective.
-Second, coreference in news introduces high–scoring entities with repeated semantics, which harms the diversity of the keyword set.
-To optimize relevance and diversity under these constraints, We propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR) as shown in Algorithm 1.
+Second, coreference in news introduces high-scoring entities with repeated semantics, which harms the diversity of the keyword set.
+To optimize relevance and diversity under these constraints, we propose an improved MMR Carbonell and Goldstein (1998) algorithm, namely Salience-Calibrated MMR (SC-MMR).
 At the $k$-th iteration, SC-MMR evaluates the score of a candidate entity $U_{i}$ by
@@
 $\mathcal{S}(U_{i},U_{j})=\frac{\mathbf{v}_{U_{i}}\cdot\mathbf{v}_{U_{j}}}{\|\mathbf{v}_{U_{i}}\|\,\|\mathbf{v}_{U_{j}}\|}.$ (4b)
-The core innovation of SC-MMR is a dynamic decaying weight $\lambda_{k}$.
-At the early stage, the weight places more emphasis on an entity’s own salience.
-As the number of selected entities $k$ increases, the weight adaptively decreases for high diversity and low redundancy.
-The computation of $\lambda_{k}$ is
-$\lambda_{k}=\max\left[0.1\,,\,1.2-\frac{\exp(0.2k-0.5)}{e}\right].$ (5)
-We further replace termination strategies that rely on a fixed score with a dynamic criterion based on the relative change of the MMR score.
+The key innovation of SC-MMR is to introduce a weight schedule $\lambda_{k}$ that changes with the number of selected keywords k, so that the focus gradually shifts from relevance to diversity.
+We adopt an annealing schedule with a lower bound:
+$\lambda_{k}=\max(\lambda_{\min},\lambda_{\max}-\exp(\alpha\cdot k-\beta)).$ (5)
+This form ensures that $\lambda_{k}$ decreases monotonically with $k$ while retaining a non-zero salience weight via $\lambda_{\min}$ to prevent the diversity term from fully dominating.
+We further introduce a dynamic termination rule based on the relative change of the MMR score.
 The iteration continues only when the next best candidate $\mathcal{U}_{k+1}^{*}$ satisfies $\text{MMR}(\mathcal{U}_{k+1}^{*})>\gamma\cdot\text{MMR}(\mathcal{U}_{k}^{*})$, where $\gamma$ is a decay factor.
-It lets the algorithm determine the stopping point according to score distribution and prevents a decline in the quality of entities caused by diminishing marginal utility.
+This design prevents the decline in entity quality caused by diminishing marginal utility.
 3.4 Information Retrieval
 This module uses the keyword set $\mathcal{K}=\{k_{1},k_{2},\ldots,k_{N}\}$ distilled in previous modules to build a comprehensive external knowledge base $\mathcal{E}$, serving as supplementary context for the subsequent detection stage.
-We implement a dual source retrieval that gathers information from the open web and Wikipedia in parallel, integrating the timeliness of the former with the authoritativeness of the latter.
-Open Web. We compose all keywords into an aggregated query for the open web $Q_{\text{web}}$ with the following logic:
-$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)\land\neg\texttt{news}\land\neg\texttt{wikipedia}.$ (6)
+We implement a dual-source retrieval that gathers information from the Open Web and Wikipedia in parallel, ensuring both timeliness and authority.
+Retrieval from Open Web. We compose all keywords into an aggregated query for the Open Web $Q_{\text{web}}$ with the following logic:
+$Q_{\text{web}}=\left(\bigwedge_{i=1}^{N}k_{i}\right)$ (6)
 The operator $\land$ requires that results relate to all keywords, which enables strict cross-keyword verification.
 We explicitly exclude news and Wikipedia sources.
-The former avoids retrieving the duplicate reports to prevent source contamination Deng et al.
+The former avoids retrieving duplicate reports to prevent source contamination Deng et al.
 (2023), and the latter prevents redundancy with the subsequent Wikipedia retrieval.
 For each returned page, we only extract its summary and search snippet as the raw corpus.
-Wikipedia.
+Retrieval from Wikipedia.
 The summary section of a Wikipedia entry usually provides the most precise definition for an entity.
@@
 After identifying the unique entry, we extract the first 3 sentences of its summary as supplementary material.
-4 Multi-LLM Interaction
-Entity-guided retrieval supplies the framework with timely external information, which together with the prior knowledge of LLMs builds a Multi-Source Information Matrix.
-This stage employs a dual-state multi-LLM interaction system that fully exploits this matrix to mitigate hallucinations Feng et al.
-(2024) and ensure accurate detection.
+4 Stage 2: Multi-LLM Interaction
+The external information provided by entity-guided retrieval and the prior knowledge of LLMs builds a Multi-Source Information Matrix.
+This stage employs a dual-state multi-LLM system that fully exploits this matrix to perform parallel and multi-perspective content reasoning and claim verification, which are finally aggregated to reach a robust and interpretable judgment.
 This stage operates in two orthogonal interaction states.
 LLM Collaboration performs parallel analyses across multiple agents to reduce inferential variance.
-LLM Debate and Judgment introduces a adversarial debate to reduce systemic bias.
-These two states form a complete reasoning chain from mining divergent evidences to making a convergent judgment.
+LLM Debate and Judgment introduces an adversarial debate to reduce systemic bias.
+These two states form a complete reasoning chain from mining divergent evidence to making a convergent judgment.
 4.1 LLM Collaboration
 LLM Collaboration state aims to reduce inferential variance.
-Through parallel analysis by multiple agents, it transforms Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate.
+Through parallel analysis by multiple agents, it transforms the Multi-Source Information Matrix from the previous stage into a structured evidence pool, which provides a stable decision basis for the subsequent adversarial debate.
 As illustrated in Figure 2, the matrix systematically integrates 4 information sources that are orthogonal and complementary in evidence distribution:
-  • In-news Information: It comes from the original text of the target news and provides the basic core content.
-  • Out-of-news Information: It comes from Open Web retrieval and offers the broadest instant facts.
-  • LLM internal Knowledge: It comes from the model’s prior knowledge and provides generalized common sense.
-  • LLM external Knowledge: It comes from Wikipedia retrieval and provides the most precise summary for the core entities.
+  • In-news Information: Comes from the original text of the target news and provides the basic core content.
+  • Out-of-news Information: Comes from Open Web retrieval and provides the most timely and broadest materials.
+  • LLM internal Knowledge: Comes from the model’s prior knowledge and provides generalized common sense.
+  • LLM external Knowledge: Comes from Wikipedia retrieval and provides the most precise summary for the core entities.
 The subsequent analysis decomposes the tasks and assigns them to agents with distinct roles, so that each agent focuses on a specific quadrant of the matrix for specialized processing.
@@
   • Grammar: Patterns of reported speech, passive voice, and negation.
-  • Emotion: Proportion of affective terms in the text and the headline, and the degree of incendiary tone.
+  • Emotion: Affective terms in the text and the headline, and the degree of incendiary tone.
   • Information Quality: Presence of clickbait, information overload, or context mismatch.
 To maintain objective independence, each dimension is evaluated in an isolated session.
-For each dimension, LLM explicitly indicates whether it reflects the news is real or fake.
-This agent uses 2 quadrants of Multi-source Information Matrix: LLM internal knowledge and in-news information from the original news text.
+For each dimension, the LLM explicitly indicates whether it reflects the news is real or fake.
+It uses 2 quadrants of the Multi-Source Information Matrix: LLM internal knowledge and in-news information.
 4.1.2 Domain-Specific Expert
-It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as “economist” or “journalist”.
-The expert with this role then analyses along the following 2 dimensions:
+It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as "economist" or "journalist".
+The expert with this role then analyzes along the following 2 dimensions:
   • Knowledge Concordance: Examine all claims, viewpoints, and details for sound reasoning; identify departures from common sense.
   • Logical Integrity: Examine argument-to-conclusion coherence with domain-specific common sense; identify logical errors or unsupported leaps.
-This agent uses 3 quadrants of Multi-source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge from Wikipedia.
+It uses 3 quadrants of the Multi-Source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge.
 4.1.3 Claim Verification
 Existing research Niu et al.
-(2024) demonstrates that claim-based fact-checking effectively serves as a reference for LLM-based detection.
+(2024) demonstrates that claim-based fact checking effectively serves as a reference for LLM-based detection.
 In our system, a serial pipeline composed of Claim Extractor and Claim Verifier deconstructs and verifies factual claims in the news text.
-These 2 agents use 3 quadrants of Multi-source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
-Claim Extractor. $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
+These 2 agents use 3 quadrants of the Multi-Source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
+Claim Extractor. The $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
 $\mathcal{M}_{\text{extractor}}(T)\rightarrow\{q_{\text{core}},\{q_{\text{sub}_{1}},\dots,q_{\text{sub}_{m}}\}\}$ (8)
@@
 This agent verifies each claim $q$ independently.
-To ensure precision and control cost Purwar et al.
-(2024), a simple retrieval augmented generation (RAG) Lewis et al.
+To ensure precision and control costs Purwar and others (2024), a simple retrieval-augmented generation (RAG) Lewis et al.
 (2020) is implemented to build a highly relevant context $\mathcal{C}_{\text{rel}}(q)$ for each $q$ from the Open Web corpus $\mathcal{I}_{\text{web}}$.
@@
 All decisions are strictly based on $\mathcal{C}_{\text{rel}}(q)$.
-The final output includes a clear label (“Supports”, “Refutes”, or “Not Enough Information”), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ for suppressing hallucinations.
+The final output includes a clear label ("Supports", "Refutes", or "Not Enough Information"), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ to suppress hallucinations.
 4.2 LLM Debate and Judgment
@@
 arguments, making the final judgment highly interpretable.
+Table 1: Accuracy (Acc.) and F1-Score (F1) comparison of few-shot / zero-shot methods on PolitiFact and GossipCop. The bold and underlined denote the best and second-best performance.
+Category Method LLM Usage PolitiFact GossipCop Accuracy F1-Score Accuracy F1-Score Few-shot PSM Nan et al.
+(2021) Non-LLM 70.00 49.15 77.44 41.73 MDFEND Nan et al.
+(2021) Non-LLM 65.50 62.30 41.27 40.20 KPL Jiang et al.
+(2022) Non-LLM 58.33 60.40 42.71 42.08 ARG Hu et al.
+(2024b) LLM-assisted 74.00 67.16 61.41 42.32 DKFND Liu et al.
+(2024) LLM-assisted 87.00 82.43 82.37 55.22 Zero-shot Auto-CoT Zhang et al.
+(2022) LLM-based 89.65 73.67 60.01 48.15 ReAct (Search API) Yao et al.
+(2023) LLM-based 74.73 67.64 74.03 47.30 HiSS Zhang and Gao (2023) LLM-based 64.82 56.80 68.81 40.40 Web Retrieval Agents Tian et al.
+(2024) LLM-based 77.83 64.88 66.62 46.54 FactAgent Li et al.
+(2024b) LLM-based 80.59 70.06 73.80 56.22 CAPE-FND Jin et al.
+(2025) LLM-based 76.81 62.43 63.79 49.17 DeepSeek-v3 DeepSeek‑AI (2024) Only-LLM 78.99 40.24 61.03 28.17 GPT-4o-mini OpenAI (2024) Only-LLM 73.58 41.66 66.73 33.18 Qwen3-32B Team (2025) Only-LLM 76.57 43.35 63.82 32.45 ZoFia (DeepSeek-v3) LLM-based 91.52 87.88 79.04 61.22 ZoFia (GPT-4o-mini) LLM-based 75.28 75.19 68.90 56.20 ZoFia (Qwen3-32B) LLM-based 84.06 81.76 69.71 59.07
 5 Experiment
 5.1 Experimental Setting
-Herein we discuss main experimental settings.
 Datasets.
@@
 political figures.
-Considering that some links in the original datasets have become invalid, we adopt subsets publicly released in Su et al.
+Considering that some links in the initial dataset have become invalid, we adopt the available version publicly re-released in Su et al.
 (2023).
-Metrics and Baselines.
-We use the accuracy (Acc.) and F1-score (F1) as the main metrics.
-To reflect performance on detecting fake news, we compute the F1-Score for the fake class.
-ZoFia compares against two groups of baselines.
-The first group includes 5 advanced Few-Shot methods: MDFEND Nan et al.
+Metrics. We use accuracy and macro F1-Score as evaluation metrics. F1-Score is less affected by data imbalance, so it serves as the primary metric for assessment.
+Baselines.
+We incorporate two groups of baselines.
+The first group includes advanced few-shot methods: PSM Ni et al.
+(2020), MDFEND Nan et al.
 (2021), ARG Hu et al.
-(2024b), PSM Ni et al.
-(2020), DKFND Liu et al.
+(2024b), DKFND Liu et al.
 (2024), and KPL Jiang et al.
-(2022), with metrics retrieved from the existing literature Jin et al.
+(2022), with reliable metrics from the existing works Jin et al.
 (2024); Hu et al.
@@
 (2024).
-The second group consists of 3 representative Zero-Shot methods that we implement for comparison: Standard Prompt, Auto-CoT Zhang et al.
-(2022), and ReAct Yao et al.
-(2023) equipped with a search API.
+The second group consists of representative zero-shot methods for comparison: Auto-CoT Zhang et al.
+(2022), ReAct Yao et al.
+(2023) equipped with a search API, HiSS Zhang and Gao (2023), Web Retrieval Agents Tian et al.
+(2024), FactAgent Li et al.
+(2024b), and CAPE-FND Jin et al.
+(2025).
+We implement ZoFia based on DeepSeek-v3 DeepSeek‑AI (2024), GPT-4o-mini OpenAI (2024) and Qwen3-32B Team (2025), and compare it with only-LLM inference.
+All other LLM-based zero-shot methods are based on DeepSeek-v3.
 Implementation details.
-We set the dynamic threshold for entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor for MMR as $\gamma=0.5$.
-The maximum entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry.
+We set the dynamic threshold of entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor of MMR as $\gamma=0.5$.
+The maximum number of entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry.
 The minimum similarity threshold for claim extraction is $\theta_{\text{sim}}=0.1$.
 The NER model is selected as dslim/bert-base-NER Tjong Kim Sang and De Meulder (2003), and the SBERT model is selected as all-MiniLM-L6-v2 Reimers and Gurevych (2019).
-All zero-shot methods we implement use DeepSeek-V3 DeepSeek‑AI (2024) with temperature set to 0.1 as the base model, and Brave API [^1] https://brave.com/search/api/ as the search API.
+For all base LLMs, we set the temperature to 0.
+Brave API [^1] https://brave.com/search/api/ is selected as Open Web retrieval API.
+To strictly prevent label information leakage, we set the search cutoff date to the day before the publication date of each sample URL and exclude dataset-associated fact-checking domains from open-web retrieval (gossipcop.com, politifact.com, and snopes.com).
 5.2 Main Experiment
-The experimental results are presented in Table 1.
-ZoFia demonstrates exceptional performance on both datasets, consistently outperforming existing zero-shot and few-shot baselines.
-On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.9%) and F1-score (89.8%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin, proving its powerful efficacy in fact-checking tasks.
-Table 1: Accuracy (Acc.) and F1-score (F1) comparison of few-shot / zero-shot methods on PolitiFact and GossipCop. The bold and underlined denote the best and second-best performance.
-Method PolitiFact GossipCop Acc.
-F1 Acc.
-F1 MDFEND 65.5 62.3 41.2 40.2 ARG 74.0 67.1 61.4 42.3 PSM 70.0 49.1 77.4 41.7 DKFND 87.0 67.1 82.4 42.3 KPL 58.3 60.4 42.7 42.1 Standard Prompt 79.0 40.2 61.0 28.1 Auto-CoT 89.7 73.7 56.7 43.3 ReAct (Search API) 74.7 67.6 74.0 47.3 ZoFia (Ours) 91.9 89.8 78.5 61.2
+ZoFia demonstrates exceptional performance on both datasets as shown in Table 1, consistently outperforming existing zero-shot and few-shot baselines.
+On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.52%) and F1-Score (87.88%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin.
 The detection task on the GossipCop dataset presents greater challenges, as its content often proves difficult to verify due to subjectivity and factual ambiguity.
-In this scenario, ZoFia achieves the highest F1-score (61.2%), surpassing all baseline models.
+In this scenario, ZoFia achieves the highest F1-Score (61.22%), surpassing all baseline models.
 Although DKFND attains slightly higher accuracy, its F1-Score is notably lower.
 In contrast, ZoFia’s superior F1-Score demonstrates more balanced and robust detection performance across both true and fake news categories while maintaining high precision.
-The experimental results clearly indicate that as a zero-shot method, ZoFia consistently outperforms the vast majority of few-shot and zero-shot models, showcasing its advancement and practical utility.
+The results clearly show that as a zero-shot method, ZoFia consistently outperforms all zero-shot baselines and even most few-shot methods.
+Compared to only-LLM inference, ZoFia provides stable and substantial performance gains, demonstrating the superiority of its architecture.
 5.3 Ablation Study
-To systematically assess the effectiveness of each component in ZoFia, we conduct extensive ablation on the GossipCop dataset.
-Figure 3: Performance comparison of ZoFia’s keyword extraction with other extraction methods.
-Effectiveness of Entity-Guided Retrieval.
-As shown in Figure 3, replacing our keyword extraction module with the classic TF-IDF Sparck Jones (1972) algorithm lowers the F1-Score by 9.1%.
-This finding emphasizes the need for our Hierarchical Salience.
-We also evaluate Sentence Salience Bullough et al.
-(2024) strategy, which retrieves a full sentence.
-The result shows an extreme imbalance: high Accuracy with low F1-Score.
-The reason is that using sentences as queries tends to retrieve pages that are highly similar to the original news.
-This source contamination Deng et al.
-(2023) induces confirmation bias in LLM agents.
-Table 2: Results of ablation study on GossipCop
-Component Acc.
-F1 ZoFia 78.5 61.2 w/o Open Web retrieval 71.7 58.1 w/o Wikipedia retrieval 77.7 59.9 w/o Open Web and Wikipedia retrieval 71.3 58.2 w/o linguist analysis 77.8 58.9 w/o expert analysis 76.8 57.3 w/o linguist and expert analyses 73.5 58.9 w/o claim verification 79.2 60.5 w/o debate 75.0 59.2
-Component Analysis of Multi-LLM Interaction.
-As shown in Table 2, removing Open Web retrieval causes a marked decline, which demonstrates the importance of timely external knowledge.
-Performance decreases further when all external knowledge sources are removed.
-For LLM collaboration, removing the linguist or expert weakens the framework.
-The expert contributes more, with a 3.9% drop in F1-Score.
-Moreover, removing both causes cumulative degradation.
-Removing the debate reduces the F1-Score by 2.0%, demonstrating its role in suppressing single-view bias.
+To assess the contribution of each component in ZoFia, we conduct ablation studies on GossipCop with DeepSeek-v3 and Qwen3-32B.
+As shown in Table 2, removing Open Web retrieval causes clear degradation in both settings, which shows that timely external evidence is essential for both base models.
+Removing all retrieval sources further lowers F1-Score to 58.26% on DeepSeek-v3 and 50.85% on Qwen3-32B.
+In LLM collaboration, removing either the linguist or the expert weakens the framework, and the expert contributes more; removing both leads to a larger cumulative drop.
+Removing debate also consistently reduces F1-Score, which shows that it helps mitigate single-perspective bias.
+The F1-Score drops by 2.06% on DeepSeek-v3 and by 5.00% on Qwen3-32B, and this larger drop suggests that the smaller model is more sensitive to the loss of this mechanism.
+Table 2: Ablation study results on GossipCop. (ACC.: Accuracy, F1: F1-Score)
+Component DeepSeek-v3 Qwen3-32B Acc.
+F1 Acc.
+F1 ZoFia 79.04 61.22 69.71 59.07 w/o Wikipedia retrieval 77.67 59.90 68.09 56.58 w/o Open Web retrieval 71.71 58.14 64.18 52.47 w/o All retrievals 71.37 58.26 62.65 50.85 w/o Linguist analysis 77.86 58.87 67.55 54.62 w/o Expert analysis 76.77 57.32 66.87 54.39 w/o All analyses 73.51 57.90 63.82 52.71 w/o Claim verification 79.23 60.50 67.48 55.17 w/o Debate 75.01 59.16 67.13 54.07
+5.4 Effectiveness of Entity-Guided Retrieval
+To verify the effectiveness of entity-guided retrieval in ZoFia, we fix the second stage and vary only how the query is constructed on GossipCop with Qwen3-32B.
+TF-IDF Sparck Jones (1972), PromptNER Ashok and Lipton (2023), and Only-LLM NER still build queries from keywords, but replace the entity acquisition mechanism, and their F1-Scores drop by 5.69%, 3.94%, and 6.33% relative to ZoFia.
+Sentence Salience Bullough et al.
+(2024), Core Claim, and OR Composition instead use sentence-level queries or relax conjunctive matching from AND ($\wedge$) to OR ($\vee$), and their F1-Scores fall by 12.92%, 10.26%, and 6.81%.
+These results show that the gain does not come from arbitrary keyword lists, and it depends on accurate entity extraction and strict entity-level constraints, which make the original conjunctive entity-guided query the most effective design.
+Table 3: Effectiveness of query construction variants in ZoFia on GossipCop.
+Method Accuracy F1-Score
+ZoFia 69.71 59.07
+ZoFia (w/o All retrievals) 62.65 50.85
+ZoFia (TF-IDF) 66.97 53.38
+ZoFia (PromptNER) 63.08 55.13
+ZoFia (Only-LLM NER) 61.69 52.74
+ZoFia (Sentence Salience) 68.64 46.15
+ZoFia (Core Claim) 66.65 48.81
+ZoFia (OR Composition) 64.89 52.26
 5.5 Sensitivity Analysis of SC-MMR
-The dynamic weight $\lambda_{k}$ is the most important parameter of SC-MMR introduced in Section 3.
-We conduct a sensitivity analysis before designing it.
-First, $\lambda_{k}$ is fixed to 0.5, indicating equal emphasis on relevance and diversity in keyword selection.
-As shown in Figure 4, we observe that when the number of selected keywords $k\leq 6$, F1-Score remains stable at a relatively high level, but once $k>6$, F1-Score drops sharply.
-This pattern indicates that too many keywords introduce noise and hurt detection performance.
+To motivate the design of the dynamic weight $\lambda_{k}$, we first conduct an experiment with the weight fixed at $\lambda=0.5$ to observe the direct impact of the keyword count $k$ on performance.
+As shown in Figure 3, the F1-Score remains stable when $k\leq 6$ but drops sharply beyond this point.
 Figure 3: The effect of the number of keywords $k$ on performance (F1-Score).
-Based on this finding, we designed $\lambda_{k}$ a decreasing function with an increasing decay rate, as shown in Eq.
-5.
-It remains high for small $k$ to prioritize relevance, reaching approximately 0.45 at $k=6$, but rapidly decaying for $k>6$.
-This design lowers the MMR score and smoothly applies the relative termination rule.
-It allows SC-MMR to adapt the size of the keyword set while avoiding performance loss due to redundancy.
+This phenomenon reveals an inflection point: when the number of keywords exceeds $6$, the risk of introducing noise and redundancy begins to outweigh the benefits of new information.
+It suggests that the strategic focus must shift from relevance to diversity near the critical inflection point of $k\approx 6$.
+Consequently, we adopt the annealing schedule form defined in Equation 5 and determine its parameters as $\alpha=0.3$ and $\beta=2.5$ to fit this downward trend.
+5.6 Efficiency of Token Utilization
+To investigate ZoFia’s reasoning efficiency, a controlled comparison experiment is conducted.
+We provide CoT Wei et al.
+(2022), FactAgent Li et al.
+(2024b), and ZoFia with the same materials and impose a unified limit on output tokens to evaluate their performance.
+Since ZoFia cannot complete full reasoning below 400 tokens, we set its evaluation range to 400 tokens and above.
+Figure 4: Performance (F1-Score) comparison of methods under maximum output token limits.
+The results are shown in Figure 4.
+At all budget points above 400 tokens, ZoFia achieves a clearly higher F1-Score than other baseline methods, which demonstrates high reasoning efficiency.
+In contrast to CoT and FactAgent, whose performance saturates after around 600 tokens, ZoFia shows a stable upward trend even at 1200 tokens.
+This indicates that ZoFia can consistently convert tokens into performance gains within an acceptable budget.
+Note that this experiment only constrains output tokens, because prompt caching Gim et al. (2024) can extensively amortize input overhead, making output cost more critical for efficiency.
 6 Conclusion
-We propose ZoFia, a training-free zero-shot framework for fake news detection, to address the fundamental conflict between the static prior knowledge of large language models (LLMs) and the uncertainty of dynamic news streams.
-ZoFia uses a two-stage pipeline to tackle this challenge.
-In the first stage, our novel Hierarchical Salience and SC-MMR algorithms extract a set of informative and low-redundancy keywords from the news text, and then build a multi-source information matrix that integrates LLM internal and external knowledge with in-news and out-of-news information.
-In the second stage, a dual-state multi-LLM interaction processes this matrix.
-The agents first perform collaborative analysis to reduce inferential variance, and then conduct adversarial debate to suppress systemic bias, finally delivering an interpretable and robust judgment.
-Comprehensive experiments on two widely used datasets show that ZoFia outperforms existing zero-shot baselines and the majority of few-shot methods.
-Detailed ablation study further demonstrates the importance and effectiveness of each core component in the framework.
+In this paper, we propose ZoFia, a retrieval-augmented multi-agent zero-shot framework for fake news detection, to address the cognitive conflict of a single LLM between content reasoning and fact checking.
+It first utilizes our novel Hierarchical Salience and Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to accurately extract core entities from news text, which then guide dual-source retrieval from Open Web and Wikipedia.
+Next, the multi-agent system conducts analysis and verification in parallel and makes a final verdict via adversarial debate.
+This process effectively reduces confirmation bias from a single reasoning perspective and ensures robust and explainable results.
+Comprehensive experiments on two public datasets show that ZoFia outperforms existing zero-shot baselines in both performance and efficiency.
 Limitations
-Though ZoFia show strong detection capacities, its application and evaluation face multiple constraints.
-Due to copyright and privacy concerns, there has been a recent lack of high-quality, and continuously updated public datasets.
+Though ZoFia shows strong detection capabilities, its application and evaluation face multiple constraints.
+Due to copyright and privacy concerns, there has been a recent lack of high-quality, continuously updated public datasets.
 It prevents us from evaluating ZoFia on the most recent news.
 Building the next generation of benchmark datasets that meet ethical standards and reflect real-world information dynamics is a critical step in this field.
-From the framework’s perspective, ZoFia depends on multiple LLMs, which introduces nontrivial computation cost and inference latency.
 The efficiency of using external retrieval can also be improved.
-Since it is not our main focus, we integrate only a lightweight retrieval augmented generation (RAG) module in claim verification.
-Future work that adopts more advanced RAG architectures, such as re-ranking model and more advanced mechanism, may further strengthen the exploitation of external knowledge.
-We plan to conduct more comprehensive experimental benchmark evaluation over the detection capabilities of LLMs and Multi-agent system for fake news Guo et al.
+Since it is not our main focus, we integrate only a lightweight retrieval-augmented generation (RAG) module for claim verification.
+Future work that adopts more advanced RAG architectures, such as a re-ranking model and more advanced mechanisms, may further strengthen the exploitation of external knowledge.
+We plan to conduct more comprehensive benchmark evaluations of the detection capabilities of LLMs and multi-agent systems for fake news Guo et al.
 (2025); Kuntur et al.
@@
 Another is to study how different modalities interact during the debate process to achieve effective fusion.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No. 62472410). We thank the reviewers and chairs for their constructive feedback.
+Appendix A Implementation Details of Modules in Entity-Guided Retrieval
+A.1 Entity Extraction
+The goal of entity extraction is to provide a stable set of semantic anchors for subsequent retrieval.
+To prevent retrieval link failures caused by entity sparsity or confidence distribution shifts, we introduce dynamic confidence threshold filtering after extraction, maintaining a controllable balance between the quality and usability of the entity set.
+Algorithm 2 details the implementation of this process.
+The inputs are the news text $\mathcal{T}$, the Named Entity Recognition (NER) model $\mathcal{M}_{\text{NER}}$, an initial threshold $\tau_{0}$, and a threshold decay step $\Delta$.
+The output is the filtered entity set $\mathcal{E}_{\text{sel}}$.
+The algorithm first runs NER on the full text to obtain the raw entity set $\mathcal{E}_{\text{raw}}=\mathcal{M}_{\text{NER}}(\mathcal{T})$, where each entity $e$ comes with a model confidence score $e.\text{ner\_score}$ that characterizes the reliability of the entity boundary and type prediction.
+If $\mathcal{E}_{\text{raw}}$ is empty, the algorithm directly returns an empty set.
+This indicates that the text lacks identifiable entity signals or the model cannot provide reliable predictions, in which case the retrieval stage should degrade to coarser-grained information clues or skip the entity-guided strategy.
+When $\mathcal{E}_{\text{raw}}$ is not empty, the algorithm employs dynamic threshold filtering to construct $\mathcal{E}_{\text{sel}}$.
+Specifically, it starts with a high initial threshold $\tau_{0}$ and retains only entities satisfying $e.\text{ner\_score}\geq\tau$, prioritizing entity quality and precision.
+If $\mathcal{E}_{\text{sel}}$ remains empty under the current threshold, the threshold decreases stepwise by $\Delta$ until the system selects at least one entity or the threshold drops to a lower bound of $0.1$.
+This mechanism reflects a clear constraint: entity extraction must provide a non-empty semantic entry point for subsequent retrieval; otherwise, retrieval degenerates into anchor-less generalized queries.
+Meanwhile, the lower bound threshold $0.1$ prevents the uncontrolled introduction of low-confidence noisy entities, thereby maintaining a baseline quality while ensuring usability.
+Dynamic threshold filtering transforms entity selection from a fixed hyperparameter setting into a sample-adaptive process.
+This enables the system to cover both entity-dense and entity-sparse news narratives, providing a stable candidate foundation for subsequent Hierarchical Salience and SC-MMR.
+Input: $\mathcal{T}$: News text;
+$\mathcal{M}_{\text{NER}}$: NER model;
+$\tau_{0}$: Initial confidence threshold;
+$\Delta$: Threshold decay step
+Output: $\mathcal{E}_{\text{sel}}$: Selected entity set
+$\mathcal{E}_{\text{raw}}\leftarrow\mathcal{M}_{\text{NER}}(\mathcal{T})$
+if $\mathcal{E}_{\text{raw}}=\emptyset$ then
+ return $\emptyset$
+$\tau\leftarrow\tau_{0}$; $\mathcal{E}_{\text{sel}}\leftarrow\emptyset$
+while $\mathcal{E}_{\text{sel}}=\emptyset\ \wedge\ \tau\geq 0.1$ do
+ $\mathcal{E}_{\text{sel}}\leftarrow\{e\in\mathcal{E}_{\text{raw}}\mid e.\text{ner\_score}\geq\tau\}$
+ $\tau\leftarrow\tau-\Delta$
+return $\mathcal{E}_{\text{sel}}$
+Algorithm 2 Entity Extraction and Dynamic-Threshold Filtering
+A.2 SC-MMR Algorithm
+A larger set of query keywords is not necessarily better, as it defines the boundary of the external evidence space.
+SC-MMR is designed to select a set of core entities that are sufficiently informative yet minimally redundant within a limited scale, which makes subsequent retrieval more stable and controllable.
+Algorithm 3 presents the complete procedure of SC-MMR.
+The input is a mapping from entities to Hierarchical Salience scores $D:\{U_{i}\mapsto S_{\text{Hier}}(U_{i})\}$, and a model $M_{\text{SBERT}}$ for calculating entity vector representations.
+The output is the final selected entity set $U_{\text{selected}}$.
+The algorithm first performs one-time pre-processing on the candidate entity set $U=\text{keys}(D)$.
+It uses $M_{\text{SBERT}}$ to encode each entity into a vector and constructs a pairwise similarity matrix $M_{\text{sim}}$, where $M_{\text{sim}}[i,j]$ measures the semantic proximity between entities $U_{i}$ and $U_{j}$.
+Pre-computing this matrix moves repetitive similarity calculations out of the iterative loop, so subsequent rounds only require table look-ups and maximum value operations.
+The algorithm then enters the selection process.
+It first selects the entity with the highest salience score $D[U_{i}]$ from the candidate set $U_{\text{candidate}}$ as the seed $U^{*}$ and adds it to $U_{\text{selected}}$.
+This initialization ensures the first step is entirely driven by relevance.
+It then enters a ‘while‘ loop, where the goal of each round is to find the entity among the remaining candidates that yields the highest combined score of relevance and diversity.
+For any candidate $U_{i}$, the algorithm first calculates:
+$\text{sim}_{\max}=\max_{U_{j}\in U_{\text{selected}}}M_{\text{sim}}[i,j],$
+which represents the similarity between the candidate and the most similar entity in the current selected set, indicating the maximum redundancy that adding $U_{i}$ might introduce.
+Subsequently, it scores the candidate using:
+$\text{MMR}_{\text{curr}}=\lambda_{k}\cdot D[U_{i}]-(1-\lambda_{k})\cdot\text{sim}_{\max}$
+Here, $\lambda_{k}$ is a weight schedule that varies with the number of selected entities $k=|U_{\text{selected}}|$, with a lower bound of 0.1.
+Intuitively, a larger $\lambda_{k}$ at small $k$ favors high-salience core entities to ensure factual coverage.
+As $k$ increases, $\lambda_{k}$ gradually decreases to strengthen the diversity penalty.
+This suppresses the repetitive inclusion of coreferential or semantically close entities, thereby mitigating query drift caused by keyword inflation.
+Each round iterates through $U_{\text{candidate}}$ to obtain the global optimal $U_{\text{best}}$ and $\text{MMR}_{\text{best}}$ as the best incremental choice.
+The algorithm does not fix the number of output keywords but adaptively stops via a relative change termination criterion.
+Let the optimal score adopted in the previous round be $\text{MMR}^{*}_{\text{prev}}$.
+If the current round’s $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}^{*}_{\text{prev}}$, the system considers the marginal gain significantly decayed.
+Since continuing to add new keywords is likely to introduce noise and redundancy, the loop terminates early.
+Otherwise, it updates $\text{MMR}^{*}_{\text{prev}}\leftarrow\text{MMR}_{\text{best}}$, adds $U_{\text{best}}$ to $U_{\text{selected}}$, and removes the entity from $U_{\text{candidate}}$.
+The finally returned $U_{\text{selected}}$ thus reflects three properties: early stages prioritize salience for key coverage, later stages use diversity penalties to suppress redundancy, and the termination rule automatically truncates the set size when marginal utility declines.
+Input: $\mathcal{D}:\{U_{i}\to\mathcal{S}_{\text{Hier}}(U_{i})\}$: Mapping from entity to its salience score;
+$\mathcal{M}_{\text{SBERT}}$: SBERT encoder for entity embeddings;
+$\gamma$: Decay factor for termination criterion
+Output: $\mathcal{U}_{\text{selected}}$: Final set of selected entities
+$\mathcal{U}\leftarrow\text{keys}(\mathcal{D})$; $\mathbf{V}\leftarrow\{\mathcal{M}_{\text{SBERT}}(U_{i})\mid\forall U_{i}\in\mathcal{U}\}$
+Compute pairwise similarity matrix $\mathbf{M}_{\text{sim}}$ from $\mathbf{V}$
+if $\mathcal{U}_{\text{candidate}}=\emptyset$ then
+ return $\mathcal{U}_{\text{selected}}$
+$U^{*}\leftarrow\mathop{\arg\max}_{U_{i}\in\mathcal{U}_{\text{candidate}}}\mathcal{D}[U_{i}]$
+$\mathcal{U}_{\text{selected}}\leftarrow\{U^{*}\}$; $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}\setminus\{U^{*}\}$
+$\text{MMR}_{\text{prev}}^{*}\leftarrow 1.0$
+while $\mathcal{U}_{\text{candidate}}\neq\emptyset$ do
+ $k\leftarrow|\mathcal{U}_{\text{selected}}|$
+ $\lambda_{k}\leftarrow\max(0.1\,,\,1.0-e^{0.3k-2.5})$
+ $\text{MMR}_{\text{best}}\leftarrow$-$\infty$; $U_{\text{best}}\leftarrow\text{null}$
+ foreach $U_{i}\in\mathcal{U}_{\text{candidate}}$ do
+    $\text{sim}_{\max}\leftarrow\max_{U_{j}\in\mathcal{U}_{\text{selected}}}\mathbf{M}_{\text{sim}}[i,j]$
+    $\text{MMR}_{\text{curr}}\leftarrow\lambda_{k}\cdot\mathcal{D}[U_{i}]-(1-\lambda_{k})\cdot\text{sim}_{\max}$
+    if $\text{MMR}_{\text{curr}}>\text{MMR}_{\text{best}}$ then
+       $\text{MMR}_{\text{best}}\leftarrow\text{MMR}_{\text{curr}}$; $U_{\text{best}}\leftarrow U_{i}$
+ if $\text{MMR}_{\text{best}}\leq\gamma\cdot\text{MMR}_{\text{prev}}^{*}$ then
+    break
+ $\text{MMR}_{\text{prev}}^{*}\leftarrow\text{MMR}_{\text{best}}$
+ $\mathcal{U}_{\text{selected}}\leftarrow\mathcal{U}_{\text{selected}}\cup\{U_{\text{best}}\}$
+ $\mathcal{U}_{\text{candidate}}\leftarrow\mathcal{U}_{\text{candidate}}\setminus\{U_{\text{best}}\}$
+return $\mathcal{U}_{\text{selected}}$
+Algorithm 3 Salience-Calibrated Maximal Marginal Relevance (SC-MMR)
+Appendix B Prompts of LLM Agents in Multi-Agent Interaction
+B.1 Prompt of Linguist
+B.2 Prompt of Expert
+B.3 Prompt of Claim Extractor
+B.4 Prompt of Claim Verifier
+B.5 Prompt of Debater
+B.6 Prompt of Judge
+If the debate reaches the maximum of 5 rounds and the judge still outputs $I$, the system terminates and returns an insufficient-evidence decision, which is counted as an incorrect case in metric computation.
+Appendix C Case Study
+This study selects a sample from the GossipCop dataset to compare two reasoning settings, ZoFia and CoT.
+Both methods use DeepSeek-v3 as the base model and share the same retrieval information from Open Web and Wikipedia.
+Figure 5 shows the source article together with truncated reasoning traces from ZoFia and CoT.
+Figure 5: Case study on a sample from the GossipCop dataset.
+The figure shows the news article of the case and the reasoning traces produced by ZoFia and CoT.
+Due to space constraints, we omit part of the reasoning with “[… …]”.
+The news claims Jamie Foxx is preparing to marry Katie Holmes, citing anonymous sources like Radar Online.
+The narrative places the wedding in Paris and adds concrete details such as wedding dress designs.
+The text follows the common style of gossip fake news.
+It uses retractable phrasings like "reportedly" and "alleged" with emotional sentences like "shook Katie’s identity".
+This approach sacrifices verifiability for dramatic effect.
+The detection challenge comes from semantic overlap.
+The couple’s relationship rumors have been widely discussed in the media.
+LLMs can easily slip from concluding that a relationship existed to believing that the marriage news is true.
+ZoFia provides a more controllable reasoning chain.
+In Entity-Guided Retrieval, Hierarchical Salience and SC-MMR narrow the keywords to Jamie Foxx and Katie Holmes.
+Dual-source retrieval simultaneously pulls web evidence and Wikipedia overviews.
+During the subsequent Multi-LLM Interaction, the task is split according to a Multi-Source Information Matrix.
+Content reasoning and fact-checking are assigned to distinct roles.
+The Linguist LLM identifies vague restatements and emotional stacking based on 5-dimensional linguistic features.
+The Expert LLM points out the reliance on anonymous sources and the logical leap of extrapolating marriage news from ambiguous statements.
+The Claim Verifier LLM rewrites the text into verifiable claims and retrieves counter-evidence for them.
+A representative for Foxx explicitly denies the engagement rumors, and authoritative reports describe their relationship as "casual", so the core claim is judged as REFUTED.
+In the debate phase, the Con.
+debater LLM consistently uses direct counter-evidence to suppress narrative-based arguments like the richness of details.
+Based on this, the Judge rules the news as FAKE.
+CoT takes an opposite path.
+The reasoning chain initially accepts the news narrative, locking in its stance early.
+It subsequently treats the sufficiency of details as a signal of truth.
+Retrieved information confirms the two had a low-profile relationship.
+The chain extrapolates this circumstantial evidence to support the marriage news, exhibiting a cognitive shortcut where retrievability replaces truthfulness.
+Counter-evidence also exists, but the denial statements about the relationship do not become the basis for the final judgment.
+Instead, the chain interprets Foxx’s statement "I’m single" as a strategy to protect his privacy.
+The conflicting information is absorbed into the narrative, leading to a final judgment of REAL.
+As mentioned in the Abstract and Introduction, LLMs are prone to early stance locking and confirmation bias, and linear CoT compresses subsequent reasoning into a rationalization of its initial judgment.
+When retrieval is introduced, the model becomes more susceptible to retrieval availability, substituting the retrievability of information for a factual judgment.
+A single LLM is also constrained by its cognitive load.
+It struggles to simultaneously handle content reasoning and external evidence verification.
+This makes it easier for counter-evidence to be rewritten into a self-consistent narrative, leading to thought degeneration.
+ZoFia addresses this by splitting tasks among multiple roles and using an adversarial debate to enforce a balance between supporting and refuting perspectives.
+The Judge terminates the process and outputs a decision when sufficient evidence is gathered, effectively breaking the erroneous reasoning chain.
+Appendix D Time Cost of the Full Pipeline
+We report the full inference latency of ZoFia by breaking it down into two parts.
+These are the LLM API call latency and the computation latency of non-LLM modules.
+All results are averaged over samples from the GossipCop dataset.
+The non-LLM modules execute on a single NVIDIA V100 GPU, and all LLM-driven modules call the GPT-4o-mini API under the same deployment setup.
+For the LLM-driven modules, we report latency in terms of relative time normalized by $T$.
+We define $T$ as the average latency of a single linguist analysis call under the same setup as shown in Table 4.
+This normalization is necessary because absolute latency varies significantly with model choice, API call method, and computation resource.
+Even in a fixed environment with the same model, the call latency is still affected by the complexity of the input and output.
+Using $T$ as a baseline mitigates these uncontrollable fluctuations and provides a more stable, comparable measure of the relative complexity among the different LLM-driven modules.
+Under our experimental setting, the average $T$ of a single GPT-4o-mini API call is approximately 2.616s.
+Table 4: LLM API call latency in relative time, normalized by $T$, where $T$ is the mean latency of one linguist analysis call measured in the same setting.
+LLM-driven module Relative time Notes
+Linguist analysis $1.00T$ per instance
+Expert triage $0.84T$ per instance
+Expert analysis $1.27T$ per instance
+Claim extraction $1.24T$ per instance
+Claim verification $1.33T$ per claim
+Debate and judgment $3.28T$ per round
+Total (with parallelism) $13.39T$ per instance
+The overall end-to-end latency also benefits from parallel execution.
+Specifically, the linguist analysis module, the expert triage plus expert analysis module, and the claim extraction plus claim verification module can run in parallel.
+This is because they operate on the same sample but depend on different intermediate signals.
+The total relative time in Table 4 already accounts for this parallel scheduling and incorporates sample-level statistics.
+The average number of claims per sample is 3.24, and the average number of debate rounds is 2.39.
+Under this setup, the end-to-end LLM-related latency after considering parallelism is $13.39T$ per sample.
+As a supplement, Table 5 presents the latency of the non-LLM modules on a V100.
+The total non-LLM latency is 0.1991 seconds, which is generally negligible compared to the LLM call latency.
+Table 5: Latency of non-LLM modules on V100.
+Non-LLM module Time (s)
+Keyword Extraction 0.1140
+RAG Load 0.0778
+RAG Retrieve 0.0073
+Total 0.1991
+In fake news detection, latency is generally not the primary constraint, as the task rarely requires a strict real-time response.
+Under this practical requirement, ZoFia’s time cost is within an acceptable range.
+The performance gains observed in the main experiments are sufficient to justify this cost.